PLEASE answer this .
You buy a commemorative coin for $25. The value of the coin increases 3.25% per year. How much will the coin be worth in 15 years? Round to the nearest cent.

Answer by MathLover1(20855)


this is geometric sequence with ;
first term, a = 25,
the coin increases 3.25%=0.0325 per year
so, common ratio, r =1+0.0325= 1.0325
the general form of a geometric sequence is
a , a r , a r ² , a r ³ , .......... , a r ⁿ
.................>>>...where a is the first term,
.................>>>...r is the common ratio,
.................>>>...nth term = a r ^ ( n - 1 )
15th term = 25 ( 1.0325 )^(15-1)
................= 25 ( 1.0325 )^14
................= 25 ( 1.5648 )
................= 39.12
after 15 years, the coin will be worth $39.10.